OpenStudy (anonymous):

What is the area of a circle with a diameter of 8? A. 4 B. 8 C. 16 D. 64

OpenStudy (bee_see):

Area of a circle= pi(r)^2. Radius is half of a diameter. so d=(1/2)(8). A=pi(4)^2=16pi

OpenStudy (anonymous):

Thanks.

OpenStudy (bee_see):

oops, one correction. I meant to write: r=(1/2)(d)=(1/2)(8)

OpenStudy (anonymous):

Oh, okay. :)